OpenStudy (anonymous):

Simplify -3xy 3 + 7xy 3 + (-xy 3) + (-3xy 3). 0 xy3 3xy3 -xy3

OpenStudy (anonymous):

We are given: x+2y=0 3x-y=3 use second equation and solve for y add y to both sides of =, then subtract 3 from both sides of = y = 3x-3 substitute for y in first equation x + 2*(3x-3) = 0 x + 6x - 6 = 0 7x = 6 x = 6/7 now substitute for x in second equation 3*(6/7) - y = 3 18/7 - y = 3 add y to both sides of = and subtract 21/7 (3) from both sides of = y = -3/7 solution set is x = 6/7 and y = -3/7 now check the answers substitute in first equation x+2y=0 6/7 + 2*(-3/7) = 0 6/7 -6/7 = 0 0 = 0 substitute in second equation 3x-y=3 3*(6/7) - (-3/7) = 3 18/7 + 3/7 = 3 21/7 = 3 3 = 3 i hope this helps